# Generation of mechanical cat-like states via optomagnomechanics

### v1

@@ III PHONON SUBTRACTION via optical pulses
In Fig. 4, we plot the Wigner function of the single/two-phonon-subtracted squeezed mechanical state. The phonon subtraction leads to the negativity of the Wigner function, demonstrating the nonclassical nature of the state. Moreover, the Wigner function exhibits a coherent superposition feature with the interference fringes shown around the origin of the phase space, which resembles that of the Schrödinger’s cat state. For this reason, it is called mechanical cat-like state and its fidelity with the cat state increases when more phonons are subtracted from the squeezed state IS20 ; Tan20 , which corresponds to a smaller successful probability and thus longer measurement time. We take the following parameters: optical wavelength $\lambda_{0}=1550$ nm, $\kappa_{c}/2\pi=3$ MHz, $g_{0}/2\pi=2$ kHz, pulse duration $t=30$ ns and power $P_{0}=50$ pW, which yield $G/2\pi=1.8$ kHz and $\tan\theta\approx 0.11$. The probability of subtracting a single phonon (two phonons) is about $\sim 1\%$ ($0.01\%$).
IV Conclusions
@@
We present a two-step scheme to prepare a cat-like state of mechanical motion in an OMM system. By applying a two-tone microwave drive field to the magnomechanical system, the mechanical mode is prepared in a squeezed thermal state in the first step. After turning off the microwave dirves, by further sending a weak red-detuned optical pulse to the optomechanical cavity, $k$ phonons can be subtracted from the squeezed mechanical mode conditioned on the detection of $k$ anti-Stokes photons from the cavity output field, which results in a mechanical cat-like state. More phonons subtracted from the squeezed state leads to a higher fidelity with the cat state, but a smaller successful probability. The work may find applications in macroscopic quantum studies and quantum sensing and in the test of collapse models.
@@ ACKNOWLEDGMENTS
This work was supported by National Key Research and Development Program of China (Grants No. 2024YFA1408900 and No. 2022YFA1405200), Zhejiang Provincial Natural Science Foundation of China (Grant No. LR25A050001), and National Natural Science Foundation of China (Grants No. 12474365 and No. 92265202).
@@
In Sec. II, we prepare a squeezed thermal mechanical state characterized by a $2\times 2$ CM $V_{b}$, which acts as the initial state $\rho_{in,b}$ for the subsequent operation of phonon subtraction. To exploit the derived propagator $U(t)$ associated with the pulse, we need to transform the form of the mechanical state from the CM to the density matrix. The squeezed thermal state can be expressed as $\rho_{in,b}=S(\xi)\rho_{th}(\bar{n})S^{\dagger}(\xi)$, where $S(\xi)$ is the squeezing operator, $\xi=re^{i\phi}$, and $\rho_{th}=\Sigma_{n}P_{n}|n\rangle\langle n|$ is the thermal state, $P_{n}=\bar{n}^{n}/(1+\bar{n})^{n+1}$, with $\bar{n}$ being the mean thermal phonon number. The characteristic parameters $r$, $\phi$, and $\bar{n}$ can be extracted from the CM $V_{b}$ via VG06
@@
with $V_{j,k}$ ($j,k=1,2$) being the elements of $V_{b}$. For the mechanical state with $V_{b}=\{\{0.045,0.14\},\{0.14,6.28\}\}$ prepared in Sec. II, we get $r=1.25$, $\phi=-0.04$, and $\bar{n}=0.013$.

### v2

@@ IV PHONON SUBTRACTION via optical pulses
In Fig. 5, we plot the Wigner function of the single/two/three-phonon-subtracted squeezed mechanical states. The phonon subtraction leads to the negativity of the Wigner function, demonstrating the nonclassical nature of the state. Moreover, the Wigner function exhibits a coherent superposition feature with the interference fringes shown around the origin of the phase space, which resembles that of the Schrödinger’s cat state. For this reason, it is called a mechanical cat-like state. We take the following parameters: optical wavelength $\lambda_{0}=1550$ nm, $\kappa_{c}/2\pi=8$ MHz, $g_{0}/2\pi=100$ Hz, pulse duration $\tau_{s}=10$ ns, and power $P_{0}=60$ nW, which yield $G/2\pi=1.8$ kHz and $\tan\theta\approx 0.11$. The probability of subtracting a single phonon, two phonons, and three phonons are on the order of $\sim 10^{-2}$, $10^{-4}$, and $10^{-6}$, respectively.
To see how close our prepared state is to the ideal cat state, we calculate the fidelity with the cat state, which is defined as
[…]
$\begin{split}{\mathcal{I}=-\frac{\pi}{2}\iint W(x,p)\left(\frac{\partial^{2}}{\partial x}+\frac{\partial^{2}}{\partial p}+2\right)W(x,p){\rm d}x{\rm d}p}.\end{split}$ (25)
The macroscopicity of the three cat-like states in Fig. 5 are 3.4, 4.7, and 5.7, respectively.
To further analyze the effects of some key parameters on our cat-like states, we plot in Figs. 6(a) and 6(b) the fidelity of the two-phonon subtracted squeezed state (to the cat state with $\alpha_{m}=$ 5.03) versus the two characteristic parameters of the squeezed thermal state, i.e., the squeezing parameter $r$ and the residual phonon number $\bar{n}$, respectively. Clearly, the fidelity increases with the squeezing when $r<1$, but a too large $r$ is not helpful as it reduces the fidelity with the cat state due to the distortion of the two superposed wave packets. The fidelity decreases with the residual phonon number $\bar{n}$, indicating that the optimal situation corresponds to the squeezed vacuum state ($\bar{n}=0$) since the ideal cat state $|\psi_{\alpha_{m}}\rangle_{\pm}$ is a pure state. In Fig. 6(c), we further plot the negativity of the Wigner function of the two-phonon subtracted squeezed state versus $\bar{n}$, which is defined as the volume of the negative Wigner distributions $\mathcal{N}=\frac{1}{2}(\iint|W(x,p)|dxdp-1)$} AK04 . Since the negativity only appears in the interference region and the Wigner becomes all non-negative when there is no interference. In this sense, the negativity can be used to infer the interference visibility of the cat-like state. Figure 6(c) shows that the interference visibility, as well as the nonclassicality, reduces with $\bar{n}$, and the optimal situation also corresponds to a vanishing $\bar{n}\to 0$.
Figure 6: The fidelity of the two-phonon subtracted squeezed thermal state to the cat state versus (a) the squeezing parameter $r$ and (b) the residual phonon number $\bar{n}$. (c) The corresponding negativity of the Wigner function versus $\bar{n}$. We fix $\bar{n}=0.014$ in (a) and $r=1.23$ in (b) and (c). The other parameters are the same as in Fig. 5.
Finally, when the mechanical motion is prepared in the cat-like state, another red-detuned optical pulse with duration $\tau_{r}$ is sent to the cavity to map the mechanical state to the cavity output field Tan20 ; Jie18a , from which the tomography of the mechanical state can be subsequently implemented.
Compared with the existing protocols using a pure optomechanical system IS20 ; Tan20 , the advantage of our protocol employing a magnomechanical system in the first step (squeezing mechanical motion) lies in the fact that the magnon dissipation rate of the YIG crystal is the lowest among all magnetic materials Chumak : $\kappa_{m}$ is around 1 MHz, which is much smaller than the mechanical frequency of both a YIG sphere Zuo and a YIG micro bridge PRAp . Thus, for the magnomechanical system, the resolved-sideband limit $\kappa_{m}\ll\omega_{b}$ is always well satisfied, which remains challenging for many optomechanical configurations MA14 . The resolved-sideband limit is required for the reservoir-engineering method to be effective and competent to generate strong mechanical squeezing. Moreover, a smaller magnon dissipation rate leads to a larger cooperativity, which is a key parameter to achieve both strong squeezing and high purity of the mechanical state IS20 .
We now analyze the effects of other experimental imperfections on our protocol. For photon detection and transmission loss, the photon loss reduces the heralding probability IS20 ; JL20 . For the anti-Stoke photon that fails to be detected (due to finite detection efficiency), it is not recorded and thus this event is discarded, so lower detection efficiency implies longer measurement time. For dark counts and false heralding, it means that the detector clicks while the anti-Stoke scattering is not actually activated. This event is recorded, but the phonon is not subtracted, and thus the mechanical mode remains in the squeezed state. This can be discriminated by subsequent tomography, where the true heralding corresponds to negative Wigner around the origin of phase space (Fig. 5), while the false heralding gives a positive Wigner at the origin (Fig. 2(b)).
V Conclusions
@@
We present a two-step pulse scheme to prepare a cat-like state of mechanical motion using an OMM system. By applying a two-tone microwave drive field to the magnomechanical system, the mechanical mode is prepared in a squeezed thermal state in the first step. By further sending a weak red-detuned optical pulse to the optomechanical cavity in the second step, $k$ phonons can be subtracted from the squeezed state conditioned on the detection of $k$ anti-Stokes photons from the cavity output field, which results in a mechanical cat-like state. More phonons subtracted from the squeezed state correspond to a higher fidelity to the cat state, but a lower successful probability. The work may find applications in macroscopic quantum studies, quantum sensing, and the test of collapse models.
@@ ACKNOWLEDGMENTS
This work was supported by Zhejiang Provincial Natural Science Foundation of China (Grant No. LR25A050001), National Natural Science Foundation of China (Grants No. 12474365 and No. 92265202), and National Key Research and Development Program of China (Grants No. 2024YFA1408900 and No. 2022YFA1405200).
@@
In Sec. III, we prepare a squeezed thermal mechanical state characterized by a $2\times 2$ CM $V_{b}$, which acts as the initial state $\rho_{in,b}$ for the subsequent operation of phonon subtraction. To exploit the derived propagator $U(t)$ associated with the pulse, we need to transform the form of the mechanical state from the CM to the density matrix. The squeezed thermal state can be expressed as $\rho_{in,b}=S(\xi)\rho_{th}(\bar{n})S^{\dagger}(\xi)$, where $S(\xi)$ is the squeezing operator, $\xi=re^{i\phi}$, and $\rho_{th}=\Sigma_{n}P_{n}|n\rangle\langle n|$ is the thermal state, $P_{n}=\bar{n}^{n}/(1+\bar{n})^{n+1}$, with $\bar{n}$ being the mean thermal phonon number. The characteristic parameters $r$, $\phi$, and $\bar{n}$ can be extracted from the CM $V_{b}$ via VG06
@@
with $V_{j,k}$ ($j,k=1,2$) being the elements of $V_{b}$. For the mechanical state with $V_{b}=\{\{0.046,0.141\},\{0.141,6.283\}\}$ prepared in Sec. III, we get $r=1.23$, $\phi=-0.05$, and $\bar{n}=0.014$.